Omar Company prepared the following contribution format income statement based on 100,000 units of sales.

Sales………………………………………………………………………………………………$3,000,000

Variable expenses…………………………………………………………………………….1,800,000

Contribution margin………………………………………………………………………….1,200,000

Fixed expenses……………………………………………………………………………………..900,000

Net operating income………………………………………………………………………… 300,000

Required:

Using the information given above answer the questions on page 226 of the text.

The following questions 4, 5, 6, 10, & 14 are changed to read as follows:

  • 4. If sales increase by 1,000 units what would be the net operating income?
  • 5. If sales decrease by 1,000 units what would be the net operating income?
  • 6. If the selling price increases by $2 per unit and the sales volume decreases by 10,000 units what would be the net operating income?
  • 10. How many units must be sold to achieve a net operating income of $350,000?
  • 14. Use the information given in the income statement above to answer this question.
Answer #1

Selling price per unit = $3,000,000 / 100,000 = $30

Variable expenses per unit = $1,800,000 / 100,000 = $18

4. Sales in units = 100,000 + 1,000 = 101,000

Net operating income = Sales - Variable costs - Fixed costs

= (101,000 * $30) - (101,000 * $18) - $900,000

= $312,000

5. Sales in units = 100,000 - 1,000 = 99,000

Net operating income = Sales - Variable costs - Fixed costs

= (99,000 * $30) - (99,000 * $18) - $900,000

= $288,000

6. Selling price per unit = $30 + $2= $32

Sales in units = 100,000 - 10,000 = 90,000

Net operating income = Sales - Variable costs - Fixed costs

= (90,000 * $32) - (90,000 * $18) - $900,000

= $360,000

10. Contribution margin ratio = (Selling price per unit - Variable expenses per unit) / Selling price per unit

= ($30 - $18) / $30

= 0.4

Units to be sold = (Fixed expenses + Desired profit) / Contribution margin ratio

= ($900,000 + $350,000) / 0.4

= $3,125,000